wiki:access/NewSun_013

CAWCR-BoM ACCESS NWP Ngamai Migration Working Group


CAWCR-BoM ACCESS-NWP Ngamai Porting Working Group Meeting Notes

Meeting 13: Wednesday 30th October 2013, 10E Meeting Room
Present: Joerg Henrichs, Ed Habjan, Wenming Lu, Jim Fraser, Chris Tingwell, Robin Bowen, Asri Sulaiman, Ilia Bermous

Apologies: Yi Xiao, Michael Naughton, Martin Dix, Joan Fernon, Zhihong Li


Agenda

  • Items from previous meeting notes
  • AOB

MARS / SAM

  • MARS-1 now running fine.
  • Full set of fields are being archived to MARS-1 from Access-G
  • Reduced set of fields from Access-R
  • Archiving of fields from Access-C to start soon
  • No status on new/additional tape drives for MARS-1
  • ACTION: Continue monitoring


AG1

  • Post processing task being turned on
    • no plots yet
  • Plan to try and make ngamai AG1 operational by next week.
    • Looks too ambitious at this stage
  • Ngamai have not been declared operational yet
    • Meeting this afternoon
  • Re Gary's plots: grand plans with lots of additional diagnostics being considered
    • work in progress
  • Chris now running development APS1 Access-G suite on ngamai


AR1

  • Elapse times have slowed slightly with increased disk i/o load on ngamai (resulting from more post-processing tasks turned on).
  • Development AR1 starting soon.
  • AG and AR to be true real-time by tomorrow (currently delayed by 4 hrs)


AC1

  • Now running routinely..
  • Wenming plotting tasks have started running from 27th Oct. Output on the Web.
  • Near real-time


APS-0 switch off

  • Slightly delayed due to NSW fires.
  • Today's 18z AG-0 is likely the very last run.
  • Joan have sent email to all stake holders.
    • No issues expected.


ATC1

  • Joan started looking at ATC1
  • Martin's "Big Domain" reconfiguration to be implemented after Xiao's back from leave


New SMS Server

  • The new virtual server is now available, but not fully set up yet.
  • Currently SMS controlling ngamai suites are running on twister.
  • FUSE (allowing real time viewing of stdout/stderr) will only be available on the new server.


Disk Space quotas


Runtime variation

  • Can now be closed after very satisfactory conclusion.


Migration of Access Applications (UIs, SVN, TRAC, DOCS )

  • Continue regular reminder to users to test the applications they use.
    • Issues have been fixed as they are discovered.
    • Email address for issues is CAWCR access ( access@… )
  • Change the group name "access" to "access.admin" to make it the same as at NCI
    • No problem expected as gid (group-id) will remain the same
    • Need to coordinate when to do it.


Compilation on Computing nodes

  • CLOSE. Successful conclusion


Small execs

  • CLOSE. No follow up needed.


CAP Program

  • Work in progress. Wenming following up with Martin.


Verify


Solar Shut down

  • Robin to continue sending weekly emails
  • Load on solar currently still high with some users saying they are finishing in the next 1-2 weeks.
  • November is critical month for switching over of NMOC's operational suites.
  • Until raijin is fully setup, some specific research work may be run on ngamai.


NWP build documentation

  • Ilia to test executable built on rajin, run on ngamai to see if identical as executables build on ngamai running on raijin.


ksh bug

  • CLOSE. No follow up required.


AOB

  • NONE.



Next Meeting


* * * 11am Wed 13th November 2013, 10E Meeting Room * * *

---> NOTE: 10 East MEETING VENUE




-> Back to Ngamai Porting index page

--> Back to Main access page


[azs, Mon 6-7/11/2013] First draft.




Last modified 6 years ago Last modified on Feb 3, 2015 4:33:57 PM